OpenStudy (anonymous):

If x+4/4 = y+7/7, then x/4 = _______. A. y-3 B. y+3 C. y/4 D. y/7

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@Jaydog477 Yeah, I'm having some severe difficulty..

hero (hero):

It's not hard if you know this rule \[\frac{a + b}{b} = \frac{a}{b} + \frac{b}{b} = \frac{a}{b} + 1\]

hero (hero):

@kalieklab I'll help you reduce the left side, then you can try the right side: \[\frac{x + 4}{4} = \frac{x}{4} + \frac{4}{4} = \frac{x}{4} + 1\]

OpenStudy (anonymous):

@Hero would the answer be B?

hero (hero):

How did you arrive at your answer? Show your reasoning or the work you did to come up with your answer.

OpenStudy (anonymous):

@Hero My sister was looking over my shoulder and suggested it's B, so I was wondering if it's correct is all (:

hero (hero):

Do you understand the steps I have demonstrated to you above?

OpenStudy (anonymous):

@Hero Yes, I do, and I really appreciate how clear you were

hero (hero):

Have you tried simplifying \(\dfrac{y + 7}{7}\) yet?

OpenStudy (anonymous):

@Hero wouldnt it just be y?

hero (hero):

@kalieklab how did you arrive at y as a result?

OpenStudy (anonymous):

@Hero I just canceled out the numbers by simplifying

hero (hero):

Can you show your work please? Thanks. Use the draw button if necessary.

OpenStudy (anonymous):

@Hero I'm really sorry, but I'm super pressed for time right now. I'm drawing it out with a pencil next to me, but I don't think I can with the pencil

OpenStudy (anonymous):

with the draw**

hero (hero):

It is best for students show their work because that's the only way I'll be able to guide you so that you know how to solve it on your own.

OpenStudy (anonymous):

@Hero I know, and I'm really sorry but I just cannot right now. Is the answer B? I really need to move on to my next question

hero (hero):

Okay, I will post the steps this time, but in the future, figure out how to use the draw button.

OpenStudy (anonymous):

@Hero thank you so much

hero (hero):

You are given: \[\frac{x + 4}{4} = \frac{y + 7}{7}\] Apply the rule \(\frac{a + b}{b} = \frac{a}{b} + \frac{b}{b}\): \[\frac{x}{4} + \frac{4}{4} = \frac{y}{7} + \frac{7}{7}\] Apply the rule \(\frac{b}{b} = 1\): \[\frac{x}{4} + 1 = \frac{y}{7} + 1\] Subtract 1 from both sides to get: \[\frac{x}{4} = \frac{y}{7}\]

OpenStudy (anonymous):

@Hero that was so perfectly explained, thank you!